# A mini-review on combinatorial solutions to the Marcus–Lushnikov irreversible aggregation

### v2

@@ VII.1 Precision of solutions
Can we study the size-dependent correctness of the solutions by numerical means? How would prediction errors scale with NN? This could be of interest for any applied research studying finite systems far from the large-size limit, as the one in Ref. Lepek_2021 .
@@ VII.3 Non-monodisperse initial conditions
To perform such re-counting, let us assume a system of NN monomeric units containing initially dimers only (thus, containing $N/2$ clusters). Such a system of “larger” initial particles will have rescaled properties: the maximum size of a cluster that could grow in the system would be $g/2$, and the total number of clusters in the system will count $k/2$. Of course, in this case, non-zero values of $\left\langle n_{s}\right\rangle$ are possible only when $s$ is even.

### v4

@@ VII.1 Precision of solutions
Can we study the size-dependent correctness of the solutions by numerical means? How would prediction errors scale with $N$? This could be of interest for any applied research studying finite systems far from the large-size limit, as the one in Ref. Lepek_2021 .
@@ VII.3 Non-monodisperse initial conditions
To perform such re-counting, let us assume a system of $N$ monomeric units containing initially dimers only (thus, containing $N/2$ clusters). Such a system of “larger” initial particles will have rescaled properties: the maximum size of a cluster that could grow in the system would be $g/2$, and the total number of clusters in the system will count $k/2$. Of course, in this case, non-zero values of $\left\langle n_{s}\right\rangle$ are possible only when $s$ is even.